OpenStudy (anonymous):

solve the equation with the restrictions of -180< x < 90 degrees.. 2 csc x =6 i have trouble doing the csc,cot and sec.. how do i plug this into my calculator csc x= 3

OpenStudy (campbell_st):

so start by solving for csc(x) = 3 then take the reciprocal csc = 1/sin so taking the reciprocal will give sin(x) = 1/3 \[x = \sin^{-1}(\frac{1}{3})\] since you value for csc is positive sin will be positive and this only occurs in the 1st quadrant... using toe information to find x is all you need.

OpenStudy (anonymous):

okay thanks!